# PIPER: Primitive-Informed Preference-based Hierarchical Reinforcement Learning via Hindsight Relabeling

### v1

@@ 5.2 Evaluation and Results
In order to analyse our design choices, we consider multiple baselines: PIPER-No-V (PIPER without primitive informed regularization, Section 4.4), RFLAT (Single-level PEBBLE (Christiano et al., 2017) implementation with target networks), HIER (vanilla hierarchical SAC implementation), HAC (Hindsight Actor Critic (Levy et al., 2018)), DAC (Discriminator Actor Critic (Kostrikov et al., 2018)), FLAT (Single-level SAC), and RAPS Dalal et al. (2021). In Figure 2, we compare the success rate performances of PIPER with hierarchical and non-hierarchical baselines. To illustrate the advantage of primitive informed regularization, we implemented PIPER-No-V, which is PIPER without primitive-informed regularization. As seen in Figure 2, though PIPER significantly outperforms PIPER-No-V baseline in pick and place and push tasks, it only slightly outperforms PIPER-No-V in the maze, hollow and kitchen environments. This illustrates that primitive-informed regularization successfully encourages the higher-level policy to generate subgoals achievable by lower-level policies.
(a) Maze navigation
(b) Pick and place
(c) Push
(d) Hollow
(e) Kitchen
(a) Maze navigation
(b) Pick and place
(c) Push
(d) Hollow
(e) Kitchen
(a) Maze navigation
(b) Pick and place
(c) Push
(d) Hollow
(e) Kitchen
(a) Maze navigation
(b) Pick and place
(c) Push
(d) Hollow
(e) Kitchen
@@ 5.3 Ablation Analysis
Here, we empirically demonstrate the importance of our each various component technique, by performing corresponding ablation study. We analyze the effect of varying the weight parameter $\alpha$ in Figure 3, and found that setting appropriate value of hyperparameter $\alpha$ is crucial for improved performance. If $\alpha$ is too small, we lose the advantages of primitive-informed regularization, leading to poor performance. In contrast, if $\alpha$ is too large, it leads to degenerate solutions. We also analyze the effect of removing the hindsight relabeling (PIPER-No-HR ablation). In Figure 4, the empirical results show that hindsight relabeling indeed leads to improved performance. Further, we remove the target networks from PIPER (PIPER-No-Target), to analyze whether using target networks stabilizes training. As seen in Figure 5, target networks significantly improve training stability.
Finally, in Figure 6, we compare PIPER against two baselines with well-behaved, hand designed dense reward functions ($(i)$ flat SAC with dense reward function (SAC-Dense), and $(ii)$ PEBBLE with actual human preferences (PEBBLE-Dense)). This comparison aims to aid the practitioners for choosing between hand-designing dense reward functions, or implementing PIPER. We find that single level SAC implementation with explicitly hand-designed and fine-tuned dense rewards (which we call SAC-Dense baseline) outperforms PIPER in easier tasks like Maze navigation and Push environments. However, we find that as the task complexity increases, the performance of SAC-Dense degrades and PIPER is able to outperform SAC-Dense in Pick and Place and kitchen environments. In the kitchen environment, SAC-Dense fails to solve the task. We believe this to is due to two reasons:
@@ 6 Discussion
Limitations: PIPER uses L2 distance between states as the informative metric, which might be hard to compute in scenarios where the subgoals and goals are high-dimensional (e.g. images). A possible way to deal with this is to find a near optimal latent representation Nachum et al. (2018a). Further, although PIPER replaces human-in-the-loop preferences via primitive-in-the-loop (PiL) approach to generate trajectory preferences, human preferences may contain additional information that PiL may be unable to capture (say, humans may prefer trajectories that are safe to traverse, which do not involve any wall collisions). A possible way to deal with this issue is to employ additional rewards in Eqn 5, which prefer a safe trajectory. Similarly, the proposed primitive regularization approach may not generate such safe subgoals in its naive form. To this end, we can additionally learn a new safe value function $V^{s}(s,g)$, which regularizes the higher agents to produce safe subgoals. Finally, the proposed target networks require additional memory and training overhead, and require setting the hyper-parameter $\tau$. However, we found in practice that the memory and training overheads are minimal, and setting $\tau$ does not cause a significant overhead. Although not explored in this work, we would like to overcome these limitations in the future.
@@
7 Impact Statement
@@ Appendix A Appendix
A.1 Implementation details
In the maze navigation task, a $7$-degree-of-freedom (7-DoF) robotic arm navigates across a four room maze, where the closed gripper (fixed at table height) has to navigate across the maze to the goal position. In pick and place task, a $7$-DoF robotic arm gripper has to navigate to the square block, pick it up and bring it to the goal position. In the push task, a $7$-DoF robotic arm gripper has to push the square block towards the goal position. In hollow task, the $7$-DOF robotic arm gripper has to pick a square hollow block and place it such that a fixed vertical structure on the table goes through the hollow block. In the kitchen task, a $9$-DoF franka robot has to pre-define a complex task to achieve the final goal. We consider the task where the robot has to open the microwave door. We also compare our approach with Discriminator Actor-Critic (Kostrikov et al., 2018), which is provided a single expert demonstration. Although not explored here, combining preference-based learning and learning from demonstrations is an interesting research avenue (Cao et al., 2020).
To ensure fair comparisons, we keep parameters including neural network layer width, number of layers, choice of optimizer, SAC implementation parameters, etc., consistent across all baselines. For implementing the RAPS baseline, we use the following lower-level behaviors: in maze navigation, we design a single primitive, reach, where the lower-level primitive travels in a straight line towards the subgoal predicted by higher level. In the pick and place, push and hollow tasks, we design three primitives: gripper-reach, where the gripper goes to given position $(x_{i},y_{i},z_{i})$; gripper-open, which opens the gripper; and gripper-close, which closes the gripper. In the kitchen task, we use the action primitives implemented in RAPS (Dalal et al., 2021).
A.1.1 Additional hyper-parameters
@@
A.2 Environment details
A.2.1 Maze navigation task
In this environment, a $7$-DOF robotic arm gripper navigates across random four room mazes. The gripper arm is kept closed and the positions of walls and gates are randomly generated. The table is discretized into a rectangular $W*H$ grid, and the vertical and horizontal wall positions $W_{P}$ and $H_{P}$ are randomly picked from $(1,W-2)$ and $(1,H-2)$ respectively. In the four room environment thus constructed, the four gate positions are randomly picked from $(1,W_{P}-1)$, $(W_{P}+1,W-2)$, $(1,H_{P}-1)$ and $(H_{P}+1,H-2)$. The height of gripper is kept fixed at table height, and it has to navigate across the maze to the goal position (red sphere).
@@ A.2.1 Maze navigation task
A.2.2 Pick and place, Push and Hollow Environments
@@
A.3 Environment visualizations
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption

### v2

@@ 5.2 Evaluation and Results
In order to analyse our design choices, we consider multiple baselines: PIPER-No-V (PIPER without primitive informed regularization, Section 4.4), RFLAT (Single-level PEBBLE (Christiano et al., 2017) implementation with target networks), HIER (vanilla hierarchical SAC implementation), HAC (Hindsight Actor Critic (Levy et al., 2018)), DAC (Discriminator Actor Critic (Kostrikov et al., 2018)), FLAT (Single-level SAC), and RAPS (Dalal et al., 2021). In Figure 2, we compare the success rate performances of PIPER with hierarchical and non-hierarchical baselines. To illustrate the advantage of primitive informed regularization, we implemented PIPER-No-V, which is PIPER without primitive-informed regularization. As seen in Figure 2, though PIPER significantly outperforms PIPER-No-V baseline in pick and place and push tasks, it only slightly outperforms PIPER-No-V in the maze, hollow and kitchen environments. This illustrates that primitive-informed regularization successfully encourages the higher-level policy to generate subgoals achievable by lower-level policies.
@@ 5.3 Ablation Analysis
Here, we empirically demonstrate the importance of our each various component technique, by performing corresponding ablation study. We analyze the effect of varying the weight parameter $\alpha$ in Figure 3, and found that setting appropriate value of hyperparameter $\alpha$ is crucial for improved performance. If $\alpha$ is too small, we lose the advantages of primitive-informed regularization, leading to poor performance. In contrast, if $\alpha$ is too large, it leads to degenerate solutions. We also analyze the effect of removing the hindsight relabeling (PIPER-No-HR ablation). In Figure 4, the empirical results show that hindsight relabeling indeed leads to improved performance. Further, we remove the target networks from PIPER (PIPER-No-Target), to analyze whether using target networks stabilizes training. As seen in Figure 5, target networks significantly improve training stability. Finally, in Figure 6, we compare PIPER against two baselines with well-behaved, hand designed dense reward functions ($(i)$ flat SAC with dense reward function (SAC-Dense), and $(ii)$ PEBBLE with actual human preferences (PEBBLE-Dense)). This comparison aims to aid the practitioners for choosing between hand-designing dense reward functions, or implementing PIPER. Although SAC-Dense outperforms PIPER in easier tasks, PIPER is able to outperform SAC-Dense in harder environments. Notably, SAC-Dense completely fails to solve the kitchen task. PEBBLE-Dense is unable to show good performance in any task. We provide further discussion in Appendix A.4.1.
@@ 6 Discussion
Limitations: PIPER uses L2 distance between states as the informative metric, which might be hard to compute in scenarios where the subgoals and goals are high-dimensional (e.g. images). A possible way to deal with this is to find a near optimal latent representation (Nachum et al., 2018a). Further, although PIPER replaces human-in-the-loop preferences via primitive-in-the-loop (PiL) approach to generate trajectory preferences, human preferences may contain additional information that PiL may be unable to capture (say, humans may prefer trajectories that are safe to traverse, which do not involve any wall collisions). A possible way to deal with this issue is to employ additional rewards in Eqn 5, which prefer a safe trajectory. Similarly, the proposed primitive regularization approach may not generate such safe subgoals in its naive form. To this end, we can additionally learn a new safe value function $V^{s}(s,g)$, which regularizes the higher agents to produce safe subgoals. Finally, the proposed target networks require additional memory and training overhead, and require setting the hyper-parameter $\tau$. However, we found in practice that the memory and training overheads are minimal, and setting $\tau$ does not cause a significant overhead. Although not explored in this work, we would like to overcome these limitations in the future.
@@
Software and Data
The implementation code and data is provided here.
Impact Statement
@@ Appendix A Appendix
A.1 Deriving the final optimum of KL-Constrained Reward Maximization Objective
In this appendix, we will derive Eqn 10 from Eqn 7. Thus, we optimize the following objective:
$\max_{\pi_{U}}\mathbb{E}_{\pi_{U}}[\sum_{t=0}^{T}(r^{s}(s,g^{*},g)-\beta\mathbb{D}_{\mathrm{KL}}[\pi^{H}(g_{t}|s_{t})\|\pi_{reg}(g_{t}|s_{t})])].$ (12)
Re-writing the above equation after expanding KL divergence:
$=\max_{\pi_{U}}\mathbb{E}_{\pi_{U}}[\sum_{t=0}^{T}(r^{s}(s,g^{*},g)-\beta\log\frac{\pi^{H}(g_{t}|s_{t})}{\pi_{reg}(g_{t}|s_{t})})]$ (13)
$=\max_{\pi_{U}}\mathbb{E}_{\pi_{U}}[\sum_{t=0}^{T}(r^{s}(s,g^{*},g)-\beta\log\pi^{H}(g_{t}|s_{t})+\beta\log\pi_{reg}(g_{t}|s_{t}))].$ (14)
Here, we substitute $\pi_{reg}$ from Eqn 8, and $m=\frac{\alpha}{\beta}$ in Equation 14,
$\begin{split}=\max_{\pi_{U}}\mathbb{E}_{\pi_{U}}[\sum_{t=0}^{T}(r^{s}(s,g^{*},g)-\beta\log\pi^{H}(g_{t}|s_{t})+\beta\log\exp(m(V_{\pi^{L}}(s_{t},g_{t})))-\beta\log\sum_{g_{t}}\exp(m(V_{\pi^{L}}(s_{t},g_{t}))))]\end{split}$ (15)
$\begin{split}=\max_{\pi_{U}}\mathbb{E}_{\pi_{U}}[\sum_{t=0}^{T}(r^{s}(s,g^{*},g)-\beta\log\pi^{H}(g_{t}|s_{t})+\alpha(V_{\pi^{L}}(s_{t},g_{t}))-\beta\log\sum_{g_{t}}\exp(m(V_{\pi^{L}}(s_{t},g_{t}))))]\end{split}$ (16)
$\begin{split}=\min_{\pi_{U}}\mathbb{E}_{\pi_{U}}[\sum_{t=0}^{T}(\log\pi^{H}(g_{t}|s_{t})-\frac{1}{\beta}(r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t})))+\log\sum_{g_{t}}\exp(m(V_{\pi^{L}}(s_{t},g_{t}))))]\end{split}$ (17)
$\begin{split}=\min_{\pi_{U}}\mathbb{E}_{\pi_{U}}[\sum_{t=0}^{T}(\log(\frac{\pi^{H}(g_{t}|s_{t})}{\exp(\frac{1}{\beta}(r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t}))))})+\log\sum_{g_{t}}\exp(m(V_{\pi^{L}}(s_{t},g_{t}))))]\end{split}$ (18)
$\begin{split}=\min_{\pi_{U}}\mathbb{E}_{\pi_{U}}[\sum_{t=0}^{T}(\log(\frac{\pi^{H}(g_{t}|s_{t})}{\frac{1}{Z(s_{t})}\exp(\frac{1}{\beta}(r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t}))))})+\log\sum_{g_{t}}\exp(m(V_{\pi^{L}}(s_{t},g_{t})))-\log Z(s_{t}))],\end{split}$ (19)
where, $Z(s_{t})=\sum_{g_{t}}\exp(\frac{1}{\beta}(r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t}))))$ is the partition function.
Note that the partition function $Z(s_{t})$ and the term $\log\sum_{g_{t}}\exp(m(V_{\pi^{L}}(s_{t},g_{t})))$ do not depend on policy $\pi_{U}$. We define
$\begin{split}\pi^{H^{*}}(g_{t}|s_{t})=\frac{1}{Z(s_{t})}\exp(\frac{1}{\beta}(r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t})))),\end{split}$ (20)
which is a valid probability distribution, as $\pi^{H^{*}}(g_{t}|s_{t})\geq 0$ and $\sum_{g_{t}}\pi^{H^{*}}(g_{t}|s_{t})=1$. We can re-organize the equation as
$\begin{split}=\min_{\pi_{U}}\mathbb{E}_{\pi_{U}}[\sum_{t=0}^{T}(\mathbb{D}_{\mathrm{KL}}[\pi^{H}(g_{t}|s_{t})\|\pi^{H^{*}}(g_{t}|s_{t})]-\log\sum_{g_{t}}\exp(m(V_{\pi^{L}}(s_{t},g_{t})))-\log Z(s_{t}))],\end{split}$ (21)
where $\pi^{H^{*}}(g_{t}|s_{t})$ is minimized when the KL-divergence is 00, according to Gibbs’ inequality. Hence, we get the optimal solution
$\begin{split}\pi^{H}(g_{t}|s_{t})=\pi^{H^{*}}(g_{t}|s_{t})=\frac{1}{Z(s_{t})}\exp(\frac{1}{\beta}(r^{s}(s,g^{*},g)+\alpha(V_{\pi^{L}}(s_{t},g_{t})))).\end{split}$ (22)
A.2 Additional Implementation details
In addition, we compare our approach with Discriminator Actor-Critic (Kostrikov et al., 2018), which is provided a single expert demonstration. Although not explored in this work, combining preference-based learning and learning from demonstrations is an interesting research avenue (Cao et al., 2020).
A.2.1 Additional hyper-parameters
@@
A.3 Environment details
A.3.1 Maze navigation task
In this environment, a $7$-DOF robotic arm gripper navigates across random four room mazes. The gripper arm is kept closed and the positions of walls and gates are randomly generated. The table is discretized into a rectangular $W*H$ grid, and the vertical and horizontal wall positions $W_{P}$ and $H_{P}$ are randomly picked from $(1,W-2)$ and $(1,H-2)$ respectively. In the four room environment thus constructed, the four gate positions are randomly picked from $(1,W_{P}-1)$, $(W_{P}+1,W-2)$, $(1,H_{P}-1)$ and $(H_{P}+1,H-2)$. The height of gripper is kept fixed at table height, and it has to navigate across the maze to the goal position(shown as red sphere).
@@ A.3.1 Maze navigation task
A.3.2 Pick and place, Push and Hollow Environments
@@
A.4 Additional experiments
In this section, we provide additional experiments, comparing PIPER with additional baselines.
A.4.1 Experiments with dense rewards baselines
We compare the performance of these baselines with PIPER in Figure 6. We find that single level SAC implementation with explicitly hand-designed and fine-tuned dense rewards (which we call SAC-Dense baseline) outperforms PIPER in easier tasks like Maze navigation and Push environments. However, we find that as the task complexity increases, the performance of SAC-Dense degrades and PIPER is able to outperform SAC-Dense in Pick and Place and kitchen environments. In the kitchen environment, SAC-Dense fails to solve the task. We believe this to is due to two reasons:
A.5 Environment visualizations